OpenStudy (anonymous):

Solve x – x3 = 2x4 + 3x2 – 5.

OpenStudy (ranga):

Left Hand side is: x - x^3 = x(1 - x^2) Right Hand side: 2x^4 + 3x^2 – 5 = 2x^4 + 5x^2 - 2x^2 - 5 = x^2(2x^2 + 5) - 1(2x^2 + 5) = (2x^2+5)(x^2-1) Equate LHS and RHS and solve for x.